OpenStudy (anonymous):

the more complete answers are really helping with my understanding, x=t-1 and y=t^2-4t i know that x rearranged gives x+1=t but i got stuck with y so far i have y+4t =t^2 can anyone help me further (eliminate parameter t and describe object)

OpenStudy (anonymous):

y+4t=t^2 solving this for t involves quadratics

OpenStudy (anonymous):

@rwinnie2 you basically use the quadratic formula \[t=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] but first u are going to get the equation in the form t^2 -4t =y then u have: a=1 b=-4 and c=0 then plug the values and find the value of t assume t is x

OpenStudy (anonymous):

wait sorry i might be wrong about all this but are u using this x+1=t to find for y+4t =t^2 ?

OpenStudy (anonymous):

if its that then u basically plug the x+1 for t in y+4t =t^2 so y+4(x+1)=(x+1)^2

OpenStudy (anonymous):

then that is y+4x+4 = (x+1)(x+1)

OpenStudy (anonymous):

(x+1)(x+1) that is x^2 +2x +1 so y+4x+4 = x^2 +2x +1 y = x^2 + 2x +1- 4x - 4 which is y=x^2 - 2x - 3

OpenStudy (anonymous):

is it making sense or am i right @rwinnie2

OpenStudy (anonymous):

ok that makes sense, but whether or not your right i have no clue, i was ill for the class and was set the classwork as homework with no additional help whatever, i did ask but he said he is busy and he will sort something out, but he wants it handed in for Thursday??????? but thank you so much i need all the help i can get